# Chernoff bound for branching random walk

### v3

@@ Abstract
Concentration inequalities, which prove to be very useful in a variety of fields, provide fairly tight bounds for large deviation probability while central limit theorem (CLT) describes the asymptotic distribution around the mean (within scope of $\sqrt{n}$ order). Harris (1963) conjectured that for a supercritical branching random walk (BRW) of i.i.d offspring and iid displacement, population’s positions in nthnth generation approach to Gaussian distribution — central limit theorem. This conjecture was latter proved by Stam (1966) and Kaplan & Asmussen (1976). Refinements and extensions followed. Yet little effort is known on large deviation probability for BRW. In this note, we propose and verify a more general and probably more formal setting of BRW. Benefiting from this definition of branching random walk, a Chernoff bound for BRW is immediately obtained. The relation between RW (random walk) and BRW is addressed.
Concentration inequality,
Chernoff bound,
@@ 1.1 Concentration inequalities
Let $(S_{i})_{i=1,2,...}$ be a random process of one-dimensional random walk on the real line. $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+...+X_{n}$; $X_{i}=S_{i}-S_{i-1}$. If some conditions are satisfied, among which independence is crucial, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT). For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-E(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [21]), also referred as Chernoff bound,
$Pr(|S_{n}-E(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
tells us how unlikely this occurrence is. In this note, by ”Chernoff bound” we refer to the tail probability inequalities of the form (1.1)—the square-exponential decay probability bound. Chernoff inequality is extended to the more general context of bounded martingale difference (say, $E(X_{n}\ |\ S_{n-1})=0$), referred as Azuma-Hoeffding inequality [4] (see [15], [28] for the survey and references therein). Refinements and extensions followed (e.g. [33], [28], [6], [7], [18], [25]). Known as concentration inequalities, these prove very useful and have a wide variety of applications in computer science, combinatorics, information theory (see e.g. [17], [3], [31], [30]).
@@ 1.2 Branching random walk (BRW)
In the literature, a branching random walk on the real line is described as follows (see, for example, [11] [8] [9]). In generation zero, an initial particle at the origin on the real line $\mathbf{R}$. It splits into a random number of child particles who form generation one. The children’s displacements, relative to their parent, correspond to a point process on $\mathbf{R}$. The children in turn split too to form the second generation, and so on. If the average split number (branching factor) is greater than one, with positive probability the number of the descendants grows exponentially through generations. Current BRW studies typically address models where the offspring’s behavior is independent of that of their previous generation (e.g. [8], [9] [10], [22], [19]). The law of large numbers and central limit theorem type results about the distribution of position are established, under some conditions of independence (for instance, i.i.d of branching and walking). Yet unlike random walk, Chernoff bound is not known so far even in the case of i.i.d aforementioned, while the minimal (and maximal) is studied by many.
Considering i.i.d. offspring (and hence independent of the parent’s position), and i.i.d. displacement, Harris [20] conjectured that the distribution of the descendants’ position of the nth generation approaches Gaussian distribution. This conjecture was proved by [32], [23]. Its extended generation-dependent versions, where offspring and displacement distribution are dependent on generation $n$, were proved by [12], [24], to mention a few. Problems where offspring and displacement are dependent of parents’ positions are studied in adhoc approaches though (e.g. [34]). Concerning the deviation from the expectation, [9] proved that for any $\delta>0$ the number of particles locating beyond $\mu-\delta$ is zero almost surely, $\mu$ denoting the scaled expectation. In another direction of estimating the rareness of the large deviation, extremum is well studied (e.g. [16] [29], [5], [13] [22], [1] [2] ). [16] showed tightness for $M_{n}-EM_{n}$, where $M_{n}$ is minimum of nthnth generation. [29] gave probability bound for the deviation of $M_{n}$, i.e. $Pr(M_{n}-Med_{n}>x)<e^{-\delta x}$, where $Med_{n}$ is the median of $M_{n}$. So far probability bound for deviation $\epsilon$ of appropriate scale ($\epsilon>0$) is established. Little is known as to how sparse the population is at a distance of $O(\sqrt{n})$ far from the mean in terms of concentration inequalities.
@@ 2 Chernoff bound for BRW
Throughout we consider BRW on $\mathbf{N}$ for notational simplicity, though the argument applies to $\mathbf{R}$. We start by a new setting for BRW, which is more general in three respects. First, in our framework both subcritical and supercritical BRW are treated equally. While traditionally subcritical BRW is considered trivial because a branching random walk process almost surely ends with zero population when $n$ is large, we see that the probability space for the survival BRW processes is well defined with infinitely large number of ancestors. Second, the underlying random walks between siblings are not assumed to be independent. Third, branching factor (birth rate) is not assumed identical across generations and siblings. The only major requirement is the independence between birth-rate and birth-place. Specifically, we define a BRW as a sequence of pairs
$\big{(}m_{i}(u),p_{i}(x\mid u)\big{)}_{i=1,2,...,n}$
where $m_{i}(u)$ is the expectation of offsprings (branching factor) of a parent at position $u$, and $p_{i}(x\mid u)$ is offsprings’ displacement pdf, the probability (or proportion) density function, or mass probability in discrete cases.
@@
$S_{n}=X_{1}+X_{2}+...+X_{n}$
be the position of a leaf in the nthnth generation, where $X_{i}$ is displacement (step size relative from the birth place) of its ithith parent. $(X_{1},X_{2},...,X_{n})$ is called a spine (a path from the origin to $S_{n}$). The central limit theorem for BRW states that almost surely a randomly chosen tree has a Gaussian “canopy”; namely, $S_{n}$ on the tree has Gaussian distribution. Let $u_{i}$ be an individual’s position, i.e. $X_{1}+...+X_{i}$, at generation $i$. At position $u_{i}$, the mean of birth rate is $m_{i}(u)$ (generally the mean of birth rate may be dependent on the birth place).
Let $M_{0}$ be the number of initial ancestors (can be infinity), and $M_{i}(u)$ the total population of ithith generation at position $u$, in the BRW forest. The population, at $i+1$ generation, produced by those $M_{i}(u)$ particles is calculated by
@@
$\sum_{x_{1}}\sum_{x_{2}}...\sum_{x_{n}}M_{0}p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})...m_{n-1}(u_{n-1})p_{n}(x_{n})$
and the law of spine $(x_{1},x_{2},...,x_{n})$
$\frac{p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})...m_{n-1}(u_{n-1})p_{n}(x_{n})}{\sum_{x_{1}}\sum_{x_{2}}...\sum_{x_{n}}p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})...m_{n-1}(u_{n-1})p_{n}(x_{n})}$ (2.1)
which can be interpreted as the proportion of $(x_{1},x_{2},...,x_{n})$ in the whole forest. As aforementioned earlier in this section, we are concerned with BRW where birth-rate is independent of birth position; i.e. $m(u)$ is not dependent on $x$. In this case, the law of spine (2.1) is reduced to
$p_{1}(x_{1})p_{2}(x_{2})...p_{n}(x_{n})$ (2.2)
which is the probability distribution of the random walk $X_{1}+X_{2}+...+X_{n}$ where $Pr(X_{i}=x\ |\ u_{i-1})=p_{i}(x)$. This observation turns questions about BRW into questions about random walk (without branching).
Remark.
@@
• A BRW of spatial homogeneity in branching (i.e. $m_{n}(u)=m_{n}$) can be studied as a random walk process (without branching).
In the following, we refer the random walk by $(p_{i})_{i=1,2,...n}$ (note this random walk is not necessarily the underlying random walk of the original BRW, though under the condition of independence between siblings’ motion, the underlying random walk is $(p_{i})$). The following theorem presents a Chernoff bound for branching random walk. An immediate corollary of this theorem establishes that, under the condition of i.i.d. which Harris (1960) considered, Chernoff bound holds.
@@
$Pr(|Q_{n}(\alpha)-na|\geq\lambda)\leq 2e^{-\frac{1}{2}c\lambda^{2}/n}$
where $na$ is the expected position111nana” is used for mean in order to be consistent with the literature of BRW, not indicating a linear relationship with $n$ for an individual in nthnth generation, $Q_{n}(\alpha)$ is $\alpha$ quantile and
$\alpha\in[e^{-\frac{1}{2}c\lambda^{2}/n},1-e^{-\frac{1}{2}c\lambda^{2}/n}]$
@@
The theorem conveys the same idea as Chernoff inequalities for random walk, stating that “almost surly” (with as small as about $e^{-O(1)\lambda^{2}/n}$ chance of exception) a BRW tree’s almost all leaves (about $1-e^{-O(1)\lambda^{2}/n}$ proportion) are concentrating around the expected position within area of $\sqrt{n}$ order (when $\lambda$ is set $O(\sqrt{n})$).
@@
$Q_{n}(1-\alpha)=\inf\{t:F(t)\geq 1-\alpha\}$
Let $\tau$ be a tree of $Z$ leaves and $p_{(\tau)}(\lambda)$ be the proportion of positions which $\geq na+\lambda$ (in the nthnth generation), i.e.
@@
By definition,
$p_{(\tau)}(\lambda)\geq\alpha\iff Q_{n}(1-\alpha)\geq na+\lambda$
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=Pr(S_{n}-na\geq\lambda)$
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\frac{\sum_{i=1}^{Z}E(\mathbf{1}_{\{z_{i}-na\geq\lambda\}})}{Z}=\frac{\sum_{1}^{Z}Pr(S_{n}-na\geq\lambda)}{Z}=Pr(S_{n}-na\geq\lambda)$ (2.3)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=Pr(S_{n}-na\geq\lambda)$
@@
$\geq\frac{1}{N_{t}}\sum_{p_{(\tau)}(\lambda)\geq\alpha}\alpha$
$=\alpha\cdot Pr(p_{(\tau)}(\lambda)\geq\alpha)$
$=\alpha\cdot Pr(Q_{n}(1-\alpha)\geq na+\lambda)$
By chernoff bound on the left-hand side of the above, for a certain constant $c>0$
$\alpha\cdot Pr(Q_{n}(1-\alpha)\geq na+\lambda)\leq e^{-c\lambda^{2}/n}$
Choosing $\alpha=e^{-\frac{1}{2}c\lambda^{2}/n}$, we have
$Pr(Q_{n}(1-e^{-\frac{1}{2}c\lambda^{2}/n})-na\geq\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.5)
and therefore
$Pr(Q_{n}(\alpha)-na\geq\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.6)
for $\alpha\leq 1-e^{-\frac{1}{2}c\lambda^{2}/n}$.
@@
$Pr(Q_{n}(\alpha)-na\leq-\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.7)
@@
Remark.
• The Chernoff bound holds for both supercritical and subcritical BRW.
• If displacements of siblings are independent of each other, with step size $X_{i}$ then, $p_{i}$ and the law of $X_{i}$ are identical (note, generally they are not equal). In other words, if $S_{n}=X_{1}+...+X_{n}$ is a random walk with $(p_{i})$ as the probability density of the increment, then the distribution of $S_{n}$ has the same ”shape” as the forest of BRW $(m_{i},p_{i})$. Bear in mind, probability space of random walk is interpreted as the forest of BRW of $(1,p_{i})$.

### v4

# Chernoff bounds for branching random walks

@@ Abstract
Concentration inequalities, which have proved very useful in a variety of fields, provide fairly tight bounds on large deviation probabilities while central limit theorem (CLT) describes the asymptotic distribution around the mean (at the $\sqrt{n}$ scale). Harris (1963) conjectured that for a supercritical branching random walk (BRW) of i.i.d offspring and i.i.d displacements, positions of individuals in nthnth generation approach to Gaussian distribution — central limit theorem. This conjecture was later proved by Stam (1966) and Kaplan & Asmussen (1976). Refinements and extensions followed. However, to the best of our knowledge, there is no corresponding existing work on concentration inequalities for BRWs. In this note, we propose a new definition of BRW, providing a more general framework. Owing to this definition, a Chernoff-type (subgaussian) bound for BRWs follows directly from the Chernoff bound for random walk. The relation between RW (random walk) and BRW is discussed.
@@ 1.1 Concentration inequalities
Let $(S_{i})_{i=1,2,...}$ be a random process of one-dimensional random walk on the real line. $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+\cdots+X_{n}$; $X_{i}=S_{i}-S_{i-1}$. Under certain conditions, among which independence is crucial, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT). For deviation beyond this neighborhood, i.e., when $|S_{n}-\mathbb{E}(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [21]), also referred to as Chernoff bound,
$\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
describes how unlikely this occurrence is. In this note, by “Chernoff bound” we refer to the tail probability inequalities of the form (1.1) – the subgaussian (square-exponential) deviation bounds. Chernoff inequalities are extended to the setting of bounded martingale difference sequences (e.g., $E(X_{n}\ |\ S_{n-1})=0$), and are then referred to as Azuma-Hoeffding inequality [4] (see [15], [29] for the survey and references therein). Refinements and extensions followed (e.g. [34], [29], [6], [7], [18], [25]). Known collectively as concentration inequalities, they prove extremely useful and have a wide variety of applications in computer science, combinatorics, information theory (see e.g. [17], [3], [32], [31]).
@@ 1.2 Branching random walk (BRW)
In the literature, a branching random walk on the real line is described as follows (see, for example, [11] [8] [9]). In generation zero, an initial particle at the origin on the real line $\mathbf{R}$. It splits into a random number of child particles who form generation one. The children’s displacements, relative to their parent, correspond to a point process on $\mathbf{R}$. The children in turn split too to form the second generation, and so on. If the average split number (branching factor) is greater than one, with positive probability the number of the descendants grows exponentially through generations. Current BRW studies typically address models where the offspring’s behavior is independent of that of their previous generation (e.g. [8], [9] [10], [22], [19]). The law of large numbers and central limit theorem type results about the distribution of position are established, under some conditions of independence (for instance, i.i.d of branching and walking). However, unlike random walk, Chernoff bound is not known so far even in the case of i.i.d aforementioned, while the minimal (and maximal) is studied by many.
Considering i.i.d. offspring (and hence independent of the parent’s position), and i.i.d. displacement, Harris [20] conjectured that the distribution of the descendants’ position of the $n$th generation approaches Gaussian distribution. This conjecture was proved by [33], [23]. Its extended generation-dependent versions, where offspring and displacement distribution are dependent on generation $n$, were proved by [12], [24], to mention a few. Problems where offspring and displacement are dependent of parents’ positions are studied in adhoc approaches though (e.g. [35]). Concerning the deviation from the expectation, [9] proved that for any $\delta>0$ the number of particles locating beyond $\mu-\delta$ is zero almost surely, $\mu$ denoting the scaled expectation. In another direction of estimating the rareness of the large deviation, extremum is well studied (e.g. [16] [30], [5], [13] [22], [1] [2] ). [16] showed tightness for $M_{n}-EM_{n}$, where $M_{n}$ is the minimal position in $n$th generation. [30] gave probability bound for the deviation of $M_{n}$, i.e. $\Pr(M_{n}-Med_{n}>x)<e^{-\delta x}$, where $Med_{n}$ is the median of $M_{n}$. While probability bounds for deviations at the $n$-scale (i.e., for fixed $\epsilon>0$ under $n^{-1}$ scaling) are by now established, little is known about how sparse the population is at distances of order $\sqrt{n}$ from the mean (position), in the concentration-inequality sense familiar from random walks.
@@ 2 Chernoff bound for BRW
Throughout we consider BRW on $\mathbf{R}$. We introduce a different definition of BRW from a new perspective, which allows us to treat BRW in a more general way in three respects: First, in our framework both subcritical and supercritical BRW are treated without distinction; while traditionally subcritical BRW is considered trivial, since a branching random walk process almost surely ends with zero population when $n$ is large, we observed that the probability space for the survival BRW processes is well defined with infinitely large number of ancestors. Second, the underlying random walks between siblings are not assumed to be independent. Third, branching factor (birth rate) is not assumed identical across generations and siblings. The only major requirement is the independence between birth-rate and birth-place.
@@
$S_{n}=X_{1}+X_{2}+\cdots+X_{n}$
be the position of a leaf in the nthnth generation, where $X_{i}$ is displacement (step size relative from the birth place) of its ithith parent. $(X_{1},X_{2},...,X_{n})$ is called a spine (a path from the origin to $S_{n}$). The central limit theorem for BRW states that almost surely a randomly chosen tree has a Gaussian “canopy”; namely, $S_{n}$ on the tree has Gaussian distribution. Let $u_{i}$ be an individual’s position, i.e. $X_{1}+\cdots+X_{i}$, at generation $i$. At position $u_{i}$, the mean of birth rate is $m_{i}(u)$ (generally the mean of birth rate may be dependent on the birth place).
In this paper, we study BRW from a new perspective. Traditionally, BRW is viewed as Galton-Watson tree weighted by spatial displacements and its study relies heavily on generating-function techniques. In contrast, we take BRW as RW ++ point process—maybe called “RWB”. In a random walk, viewed as a special BRW, produces exactly one child and is replaced by it. The intensity measure of the displacement point process $\{\xi_{1}\}$ is
$\mathbb{E}\left[\mathbf{1}_{(x,x+dx)}(\xi_{1})\right]:={d}P(x)=p(x\mid u)\,dx$
where $u$ is the parent’s position. In a general BRW, offspring form a point process with a random number of offspring $\mathcal{N}$. By Wald’s lemma [36], the corresponding intensity measure of the point process becomes
$dZ(x\mid u):=\mathbb{E}\Big[\sum_{j=1}^{\mathcal{N}}\mathbf{1}_{(x,x+dx)}(\xi)\Big]=\mathbb{E}[\mathcal{N}]\cdot\mathbb{E}\left[\mathbf{1}_{(x,x+dx)}(\xi)\right]=m(u)\cdot{d}P(x\mid u)$
or equivalently
$Z(X<x\mid u):=\mathbb{E}\Big[\sum_{j=1}^{\mathcal{N}}\mathbf{1}_{(-\infty,x)}(\xi)\Big]=\mathbb{E}[\mathcal{N}]\cdot\mathbb{E}\left[\mathbf{1}_{(-\infty,x)}(\xi)\right]=m(u)\cdot P(X<x\mid u)$
where $dP(x|u)=p(x|u)dx$ is the displacement law of a single offspring. The second equality above holds due to the independence of $\mathcal{N}$ and the offspring displacements. Here $m(u)=\mathbb{E}[\mathcal{N}]$ is the expected offspring (branching factor) by a parent at position $u$. In the above definition, the dropped subscript “1” should be replaced by the generation index $i$, which we omit for notational simplicity; in a general BRW, branching factor and the law of the offspring point process may vary across generations.
This measure $dZ(x\mid u)$ at each generation uniquely determines the law of the BRW. Specifically, the probability law of the spine (a single lineage of descendants) is
$\frac{\displaystyle\prod_{i=1}^{n}m_{i}(u_{i-1})p_{i}(x_{i}\mid u_{i-1})}{\displaystyle\int\prod_{i=1}^{n}m_{i}(u_{i-1})p_{i}(x_{i}\mid u_{i-1})\,dx_{1}dx_{2}\cdots dx_{n}}\,dx_{1}dx_{2}\cdots dx_{n}$ (2.1)
which can be interpreted as the proportion of spines $(x_{1},x_{2},...,x_{n})$ in the whole forest.
In view of this, a BRW can be defined by a sequence of pairs
$\Big(m_{i}(u),\,p_{i}(x\mid u)\Big)_{i=1,2,...}$
where $m_{i}(u)$ is the expectation of offspring (branching factor) of a parent at position $u$, and $p_{i}(x\mid u)$ is pdf of offspring’s displacement, the probability (or proportion) density function, or, in discrete cases, mass probability. This definition covers a more general cases than the traditional setting; for instance, the parents’ lifetimes need not be specified; whether they die or live, all such information is encoded in $m(u)$.
As mentioned earlier in this section, we are concerned with BRWs in which birth-rate is independent of birth position; that is, $m(u)$ does not depend on $x$. In this case, the law of spine (2.1) reduces to
$p_{1}(x_{1})p_{2}(x_{2})\cdots p_{n}(x_{n})\,dx_{1}\,dx_{2}\cdots dx_{n}$ (2.2)
which is precisely the probability measure (law) for the random walk $(X_{1},X_{2},...,X_{n})$ where
$p_{i}(x)\equiv p(x|u_{i-1})$
This observation turns questions about BRW into questions about a random walk (without branching).
To aid intuition, we illustrate the new BRW framework in the discrete case. Let $M_{0}$ be the number of initial ancestors (can be infinity), and $M_{i}(u)$ the total population of ithith generation at position $u$, in the BRW forest. The population, at $i+1$ generation, produced by those $M_{i}(u)$ particles is calculated by
@@
$\sum_{x_{1}}\sum_{x_{2}}\cdots\sum_{x_{n}}M_{0}p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})\cdots m_{n-1}(u_{n-1})p_{n}(x_{n})$
and the law of the spine $(x_{1},x_{2},...,x_{n})$
$\frac{p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})\cdots m_{n-1}(u_{n-1})p_{n}(x_{n})}{\sum_{x_{1}}\sum_{x_{2}}\cdots\sum_{x_{n}}p_{1}(x_{1})m_{1}(u_{1})p_{2}(x_{2})\cdots m_{n-1}(u_{n-1})p_{n}(x_{n})}$ (2.3)
which is the proportion of spines $(x_{1},x_{2},...,x_{n})$ in the whole forest. If the birth-rate is independent of birth position, $m(u)$ does not depend on $x$. In this case, the law of the spine (2.3) reduces to
$p_{1}(x_{1})p_{2}(x_{2})\cdots p_{n}(x_{n})$ (2.4)
which is the probability measure of the random walk $(X_{1},X_{2},...,X_{n})$ where
$p_{i}(x)\equiv\Pr(X_{i}=x\ |\ u_{i-1})$
Remarks.
@@
• A BRW of spatial homogeneity in branching (i.e. $m_{n}(u)=m_{n}$) can be studied as a random walk process (without branching). In particular,
$\int_{\left|\,S_{n}-\mathbb{E}(S_{n})\,\right|\geq\lambda}\frac{\prod_{i=1}^{n}m_{i}(u_{i-1})p_{i}(x_{i}\mid u_{i-1})}{\int\prod_{i=1}^{n}m_{i}(u_{i-1})p_{i}(x_{i}\mid u_{i-1})\,dx_{1}dx_{2}\cdots dx_{n}}\,dx_{1}dx_{2}\cdots dx_{n}$
reduces to a large deviation problem for the underlying random walk, namely
$\int_{\left|\,S_{n}-\mathbb{E}(S_{n})\,\right|\geq\lambda}\prod_{i=1}^{n}p_{i}(x_{i}\mid u_{i-1})\,dx_{1}dx_{2}\cdots dx_{n}=\Pr\left(|S_{n}-\mathbb{E}(S_{n})\geq\lambda|\right)$
In the following, we denote the random walk by $(p_{i})_{i=1,2,...n}$ (note this random walk is not necessarily the underlying random walk of the original BRW, though under the assumption of independence between siblings’ motion, the underlying random walk is $(p_{i})$). The following theorem presents a Chernoff bound for branching random walks; the classical i.i.d. model of Harris (1960) arises as a special case.
@@
$\Pr\big(|Q_{n}(\alpha)-na|\geq\lambda\big)\leq 2e^{-\frac{1}{2}c\lambda^{2}/n}$
where $na$ is the expected position111nana” is used for “mean” in order to be consistent with the literature of BRW, not indicating a linear relationship with $n$ for an individual in nthnth generation, $Q_{n}(\alpha)$ is $\alpha$ quantile and
$\alpha\in\left[e^{-\frac{1}{2}c\lambda^{2}/n},1-e^{-\frac{1}{2}c\lambda^{2}/n}\right]$
@@
The theorem conveys the same spirit as the Chernoff inequalities for random walks, stating, roughly, that with high probability (with exception probability of order $e^{-O(1)\lambda^{2}/n}$), at least a proportion $1-e^{-O(1)\lambda^{2}/n}$ of the leaves of a BRW tree are concentrated within a distance of order $\sqrt{n}$ from the expectation, when $\lambda\gtrsim\sqrt{n}$.
@@
$Q_{n}(1-\alpha)=\inf\{t:\frac{\sum_{i}\mathbf{1}_{\{z_{i}\leq t\}}}{Z^{(n)}}\geq 1-\alpha\}$
Let $\tau$ be a tree of $Z$ leaves and $p_{(\tau)}(\lambda)$ be the proportion of positions which $\geq na+\lambda$ (in the $n$th generation), i.e.
@@
By definitions,
$p_{(\tau)}(\lambda)\geq\alpha\iff Q_{n}(1-\alpha)\geq na+\lambda$ (2.5)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-na\geq\lambda)$
@@
$\mathbf{E}\left[p_{(\tau)}(\lambda)\right]=\frac{\sum_{i=1}^{Z}\mathbb{E}(\mathbf{1}_{\{z_{i}-na\geq\lambda\}})}{Z}=\frac{\sum_{1}^{Z}\Pr(S_{n}-na\geq\lambda)}{Z}=\Pr(S_{n}-na\geq\lambda)$ (2.6)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-na\geq\lambda)$
@@
$\geq\frac{1}{N_{t}}\sum_{p_{(\tau)}(\lambda)\geq\alpha}\alpha=\alpha\cdot\Pr\Big(p_{(\tau)}(\lambda)\geq\alpha\Big)$
[…]
By the well known Chernoff bound for the left-hand side of the above, it follows,
$\alpha\cdot\Pr\Big(Q_{n}(1-\alpha)\geq na+\lambda\Big)\leq e^{-c\lambda^{2}/n}$
where $c>0$ is a constant that may depend on $n$. Choosing $\alpha=e^{-\frac{1}{2}c\lambda^{2}/n}$, we have
$\Pr\Big(Q_{n}(1-e^{-\frac{1}{2}c\lambda^{2}/n})-na\geq\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.8)
[…]
$\Pr(Q_{n}(\alpha)-na\geq\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}\,;$ (2.9)
@@
$\Pr\Big(Q_{n}(\alpha)-na\leq-\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.10)
@@
Remarks.
• The Chernoff bound holds for both supercritical and subcritical BRW, because it is independent of the branching factor $m_{i}$.
• The Chernoff bound presented in this paper is a special case of the more general concentration inequalities ([28]), which holds in a setting where (a) $m_{i}(u)$ is position-dependent (and thus path-dependent), (b) neither independence nor martingale conditions are assumed for the underlying random walk processes, and (c) the boundedness assumption on the step increments $X_{i}$ is relaxed.
• If displacements of siblings are independent of each other, with step size $X_{i}$ then, $p_{i}$ and the law of $X_{i}$ are identical (note, generally they are not equal). In other words, if $S_{n}=X_{1}+\cdots+X_{n}$ is a random walk with $(p_{i})$ as the probability density of the increment, then the distribution of $S_{n}$ has the same ”shape” as the forest of BRW $(m_{i},p_{i})$. Bear in mind, probability space of random walk is interpreted as the forest of BRW of $(1,p_{i})$.